# A substellar flyby that shaped the orbits of the giant planets

### v1

@@ Abstract
The modestly eccentric and non-coplanar orbits of the giant planets pose a challenge to solar system formation theories which generally indicate that the giant planets emerged from the protoplanetary disk in nearly perfectly circular and coplanar orbits. We demonstrate that a single encounter with a 2–50 Jupiter-mass object, passing through the solar system at a perihelion distance less than 20 AU and a hyperbolic excess velocity less than $6\,\mathrm{km\,s}^{-1}$, can excite the giant planets’ eccentricities and mutual inclinations to values comparable to those observed. We estimate that there is about a 1-in-100 chance that such a flyby produces a dynamical architecture similar to that of the solar system. We describe a metric to evaluate how closely a simulated system matches the eccentricity and inclination secular modes of the solar system. The scenario of a close encounter with a substellar object offers a plausible explanation for the origin of the moderate eccentricities and inclinations and the secular architecture of the planets.
@@ 1 Introduction
Since antiquity, astronomers have sought to unravel the intricate patterns of planetary motion, refining their models over centuries to better align with empirical observations. In some ways, the historical expectations that planets move in perfect circles align with modern planet formation theories which suggest that planets emerge from protoplanetary disks in nearly circular and coplanar orbits (Goldreich and Tremaine, 1980; Artymowicz, 1992; Pollack et al., 1996; Lubow et al., 1999; Papaloizou et al., 2001; Ida and Lin, 2004; Kley and Nelson, 2012; Armitage, 2020). Consequently, the origin of the modestly eccentric and modestly non-coplanar orbits of the giant planets of the solar system has been an abiding puzzle (Tsiganis et al., 2005; Morbidelli et al., 2007; Nesvornỳ and Morbidelli, 2012; Raymond et al., 2020; Nesvorný, 2018; Clement et al., 2021; Griveaud et al., 2024). In this paper, we investigate external perturbations from flybys of stellar and substellar objects as a mechanism for the origins of the dynamical architecture of the planets.
@@
As background to our investigation, we note that the present-day eccentricities and inclinations of the solar system planets are the instantaneous manifestations of their long-term variations under mutual planetary perturbations. In the linear approximation of the Laplace-Lagrange secular perturbation theory, these long-term (secular) variations of the eccentricities are separable from those of the inclinations, and both are approximated by the superposition of linear modes whose frequencies are determined by the planetary masses and semi-major axes (Brouwer and Clemence, 1961; Murray and Dermott, 1999; Tremaine, 2023). The secular modes and their long-term variations have been the subject of many studies as they are crucial for understanding the long-term dynamical stability of the solar system (Laskar, 1988, 1990, 1996; Ito and Tanikawa, 2002; Laskar and Gastineau, 2009; Zeebe, 2015; Mogavero and Laskar, 2021, 2022; Hoang et al., 2022) as well as the climate cycles of Earth (e.g. Laskar et al., 2004, 2011). These studies show that the semi-major axes and the secular mode amplitudes of the planets are largely well preserved over billion-year timescales. Assuming that the giant planets’ masses and orbital separations were largely determined by formation processes in the primordial protoplanetary disk (with only minor changes subsequently), one infers that the amplitudes of the eccentricity and inclination secular modes must originate from some form of internal or external perturbation in the ancient solar system, after planet formation. Possible internal perturbations include planet-planet interactions, either from scattering encounters or mean motion resonances; external perturbations include stellar flybys.
@@
In our investigation of perturbations from flybys of stellar and substellar objects, we initially focus on the giant planets because their secular modes are nearly unaffected by the terrestrial planets (Laskar, 1990). Later we also include the terrestrial planets in our analysis. Previous studies to constrain possible signatures of stellar flyby events in the solar system’s history have looked at the orbital configuration of the small bodies in the outer solar system (the Kuiper belt, Sednoids, and Oort Cloud) (Levison et al., 2004; Morbidelli and Levison, 2004; Kenyon and Bromley, 2004; Brasser et al., 2006; Adams, 2010; Huang and Gladman, 2024; Pfalzner et al., 2024), and the long-term stability of the solar system (Li and Adams, 2015; Li et al., 2019; Brown and Rein, 2022; Raymond et al., 2024). Indeed, significant effort has been undertaken to investigate the stability and dynamics of planetary systems in general within the context of stellar flyby encounters (Laughlin and Adams, 1998, 2000; Malmberg et al., 2007; Spurzem et al., 2009; Malmberg et al., 2011; Hao et al., 2013; Cai et al., 2017; Li et al., 2020; Schoettler and Owen, 2024, and references therein). However, none of these investigations consider both extremely close (periastron distance less than $20\,\mathrm{AU}$) and substellar-mass objects (masses less than $50\,\mathrm{M}_{\mathrm{J}}$) together in the Sun’s primordial star cluster.
@@
In this paper, we consider close encounters from a typical open star cluster environment using an initial mass function that includes sub-stellar objects with masses ranging between $10^{-3}\,\mathrm{M}_{\odot}$ and $10^{2}\,\mathrm{M}_{\odot}$. We focus on encounters with perihelia less than $20\,\mathrm{AU}$ of the Sun. Due to the closeness of the flybys, we run full $N$-body simulations rather than analytical estimates to accurately resolve the close encounters. In total, we simulate $5\cdot 10^{4}$ flybys with the solar system giant planets initially on circular, co-planar orbits at their current semi-major axes.
@@ 2.1 Stellar Environment
We use a Monte Carlo approach to investigate the impact of various flybys on the solar system by simulating $5\cdot 10^{4}$ flyby objects from an open star cluster environment with a mass density of $n=100\,\mathrm{M_{\odot}\,pc^{-3}}$. For each flyby, we sample a mass, velocity, and impact parameter along with a random orientation centered on the Sun (Zink et al., 2020). The initial mass function (IMF) we use to generate the flyby masses $m_{\star}$ is the Maschberger IMF (Maschberger, 2013) which smoothly combines a log-normal IMF for single stars (Chabrier, 2003) with masses less than one solar mass and the standard power-law IMF (Salpeter, 1955) for masses above one solar mass. Our full IMF contains objects with masses ranging from $10^{-3}\,\mathrm{M}_{\odot}$ to $10^{2}\,\mathrm{M}_{\odot}$ and a median object mass of about $80\,\mathrm{M}_{\mathrm{J}}$. We note that extending of the stellar IMF well into the substellar regime is somewhat uncertain but is consistent with microlensing observations (Chabrier and Lenoble, 2023). We sample hyperbolic excess velocities $v_{\infty}$ from a Maxwell-Boltzmann distribution with a velocity dispersion of $1\,\mathrm{km\,s}^{-1}$ (Girard et al., 1989; Adams, 2010). We sample impact parameters $b_{\star}$ to uniformly cover the cross-sectional area out to $75\,\mathrm{AU}$. Gravitational focusing (Binney and Tremaine, 2008) naturally gives a perihelion which is less than the impact parameter with the host star. For our simulations, all perihelia come within $55\,\mathrm{AU}$ with $90\%$ coming within $20\,\mathrm{AU}$. We sample the incident angles isotropically to allow for all possible encounter geometries. We initialize the solar system with the four giant planets on circular and coplanar orbits at their current semi-major axes, and randomly initialize their true anomalies.
@@ 2.3 Metric for secular architectures
For the purpose of comparing the similarity of two systems, the approximately conserved quantity of the angular momentum deficit (AMD) (Laskar, 1997, 2000; Turrini et al., 2020) is not a good measure because the AMD can be dominated by a single planet—a common result of flyby encounters. A comparison of the instantaneous eccentricities and inclinations alone is insufficient when comparing the dynamical architectures of two systems because these represent only a snapshot of the time evolution of a system and not the overall structure. Even measuring the minimum-to-maximum range of the eccentricities and inclinations of each of the planets is insufficient because it does not measure the distribution of power spread across different secular modes. Measuring the closeness of the secular modes using their frequencies and amplitudes is also not straightforward because of the many degrees of freedom. We also would like to include the long-term variations of the secular modes, since secular modes are merely the linearized approximation of an inherently non-linear system. All the above reasons motivate us to develop a new metric which measures the distance of a simulated system from the real solar system without making any initial assumptions regarding their dynamical structure or architecture.
Let us start by defining the root-mean-square log-spectral distance (Gray and Markel, 1976; Rabiner and Juang, 1993) which compares the power spectra of two signals to evaluate the similarity of the amplitudes for each frequency as
@@
where $e_{j}$ is the eccentricity, $i_{j}$ is the inclination with respect to the invariant plane, $\Omega_{j}$ is the longitude of the ascending node, $\varpi_{j}=\omega_{j}+\Omega_{j}$ is the longitude of the periapsis, $\omega_{j}$ is the argument of the periapsis, and $\mathtt{i}=\sqrt{-1}$. We sample these values on evenly spaced time intervals over 20 Myr integrations with $N=2^{11}$ samples to then process with an FFT. We use $\mathcal{D}$ in equation (1) to build a metric that evenly weighs the signals of the complex eccentricities and complex inclinations for each planet. We define our metric as
@@
2.4 Reference Systems
In order to make a fair comparison between the flyby systems and the observed solar system, we use a reference set of solar system realizations from an ensemble of integrations (Brown and Rein, 2020). This ensures that our reference set representing the solar system’s secular architecture samples not just one single 20 Myr realization but also samples the effect of uncertainties of the initial conditions and of the small changes in the solar system’s secular modes over long times, both of which arise from nonlinear effects. Indeed, the secular mode frequencies of the giant planets change by at most $15\,\mathrm{mas\,yr}^{-1}$ and their amplitudes vary by at most $9.5\cdot 10^{-4}$ over billions of years.
We start with the nominal initial conditions that are identical to what we see today using NASA’s Jet Propulsion Laboratory (JPL) Horizons data at the J2000 epoch. Our solar system ensemble uses these nominal initial conditions that are identical in every way except for a small variation ($<0.5\,\mathrm{m}$) in the semi-major axis of Mercury from the present-day value (see Brown and Rein 2020 for more details). We use all $96$ of the 5 Gyr solar system simulations and divide them into 20 Myr non-overlapping segments from which we sample the complex eccentricities and inclinations to process with our metric.
Using the metric $\mathcal{M}$ and this set of 20 Myr segments from our reference solar system ensemble, we cross compare each segment of a random subset of $2500$ segments against another randomly selected subset of $2500$ segments to compute the distribution of values. This distribution of values of $\mathcal{M}$ quantifies how similar the solar system is to itself at different times and with slightly different initial conditions (within observational uncertainties). The range of values given by $\mathcal{M}$ for this cross comparison of the solar system are between 00 and $0.434$.
@@ 2.4 Reference Systems
We found that this procedure provides a clear way of evaluating how close two systems are to each other and, most importantly, allows us to sort and rank them in closeness. More over, this procedure avoids over-fitting to a specific manifestation of the solar system’s long and diverse dynamical evolution. What is not immediately clear is how to define the cutoff that determines if a given system is considered to be a close match to the solar system or not. For the present study, we define any flyby system with a metric value that overlaps with the reference solar system distribution as a close match to the solar system. However, we acknowledge that this choice is somewhat arbitrary and comes with several caveats. Note that the maximum value of $\mathcal{M}$ could be limited by the ensemble size. Our ensemble size is large, $2500^{2}>6\cdot 10^{6}$, so we don’t expect this to be important in practice. However, there exists the rare possibility that a solar system trajectory will eventually go unstable, for example with a Mercury–Venus collision (Laskar et al., 2004). The closer we get to such an instability, the more the trajectory diverges from what is a more ‘typical’ realization of the solar system. Such a divergent trajectory will cause the range of metric values to increase significantly above the typical range and thus lead to more closely matching flyby simulations. However, if there aren’t any significantly divergent trajectories (as is the case with our solar system ensemble, see Brown and Rein 2020), then to use anything less than the entire range of the metric would be to claim that the solar system is not a close match to itself. In summary, although it may be possible to find a better metric, the definition of what is considered a close match to the solar system will likely always be subject to some reasonable but arbitrary quantitative choices.

### v2

@@ Abstract
The modestly eccentric and non-coplanar orbits of the giant planets pose a challenge to solar system formation theories which generally indicate that the giant planets emerged from the protoplanetary disk in nearly perfectly circular and coplanar orbits. We demonstrate that a single encounter with a 2–50 Jupiter-mass object, passing through the solar system at a perihelion distance less than 20 AU and a hyperbolic excess velocity of 1–3 $\mathrm{km\,s}^{-1}$, can excite the giant planets’ eccentricities and mutual inclinations to values comparable to those observed. We describe a metric to evaluate how closely a simulated flyby system matches the eccentricity and inclination secular modes of the solar system. We estimate that there is about a 1-in-9000 chance that such a flyby occurs during the solar system’s residence in its primordial cluster and produces a dynamical architecture similar to that of the solar system. The scenario of an ancient close encounter with a substellar object offers a plausible explanation for the origin of the moderate eccentricities and inclinations and the secular architecture of the planets. We discuss some broader implications of disruptive flyby encounters on planetary systems in the Galaxy.
@@ 1 Introduction
Since antiquity, astronomers have sought to unravel the intricate patterns of planetary motion, refining their models over centuries to better align with empirical observations. In some ways, the historical expectations that planets move in perfect circles align with modern planet formation theories which suggest that planets emerge from protoplanetary disks in nearly circular and coplanar orbits (Goldreich and Tremaine, 1980; Artymowicz, 1992; Pollack et al., 1996; Lubow et al., 1999; Papaloizou et al., 2001; Ida and Lin, 2004; Kley and Nelson, 2012; Armitage, 2020). Consequently, the origin of the modestly eccentric and modestly non-coplanar orbits of the giant planets of the solar system has been an abiding puzzle (Tsiganis et al., 2005; Morbidelli et al., 2007; Nesvornỳ and Morbidelli, 2012; Raymond et al., 2020; Nesvorný, 2018; Clement et al., 2021; Griveaud et al., 2024). In this paper, we investigate external perturbations from flybys of stellar and substellar objects as a mechanism for the origins of the eccentricities and inclinations of the planets.
@@
As background to our investigation, we note that the present-day eccentricities and inclinations of the solar system planets are the instantaneous manifestations of their long-term variations under mutual planetary perturbations. In the linear approximation of the Laplace-Lagrange secular perturbation theory, these long-term (secular) variations of the eccentricities are separable from those of the inclinations, and both are approximated by the superposition of linear modes whose frequencies are determined by the planetary masses and semi-major axes (Brouwer and Clemence, 1961; Murray and Dermott, 1999; Tremaine, 2023). Henceforth, we will refer to these linear modes as the fundamental secular modes. The fundamental secular modes and their long-term variations have been the subject of many studies as they are crucial for understanding the long-term dynamical stability of the solar system (Laskar, 1988, 1990, 1996; Ito and Tanikawa, 2002; Laskar and Gastineau, 2009; Zeebe, 2015; Mogavero and Laskar, 2021, 2022; Hoang et al., 2022) as well as the climate cycles of Earth (e.g. Laskar et al., 2004, 2011). These studies show that the semi-major axes and the fundamental secular mode amplitudes of the planets are largely well preserved over billion-year timescales. Assuming that the giant planets’ masses and orbital separations were largely determined by formation processes in the primordial protoplanetary disk (with only minor changes subsequently), one infers that the amplitudes of the eccentricity and inclination secular modes must originate from some form of internal or external perturbation in the ancient solar system, after planet formation. Possible internal perturbations include planet-planet interactions, either from scattering encounters or mean motion resonances; external perturbations include stellar flybys.
@@
In our investigation of perturbations from flybys of stellar and substellar objects, we initially focus on the giant planets because their orbital evolution is nearly unaffected by the terrestrial planets (Laskar, 1990). Later we also include the terrestrial planets in our analysis. Previous studies to constrain possible signatures of stellar flyby events in the solar system’s history have looked at the orbital configuration of the small bodies in the outer solar system (the Kuiper belt, Sednoids, and Oort Cloud) (Levison et al., 2004; Morbidelli and Levison, 2004; Kenyon and Bromley, 2004; Brasser et al., 2006; Adams, 2010; Huang and Gladman, 2024; Pfalzner et al., 2024), and the long-term stability of the solar system (Li and Adams, 2015; Li et al., 2019; Brown and Rein, 2022; Raymond et al., 2024; Kaib and Raymond, 2025). Indeed, significant effort has been undertaken to investigate the stability and dynamics of planetary systems in general within the context of stellar flyby encounters (Laughlin and Adams, 1998, 2000; Malmberg et al., 2007; Spurzem et al., 2009; Malmberg et al., 2011; Hao et al., 2013; Cai et al., 2017; Li et al., 2020; Schoettler and Owen, 2024, and references therein). However, none of these investigations consider both extremely close (periastron distance less than $20\,\mathrm{AU}$) and substellar-mass objects (masses less than $50\,\mathrm{M}_{\mathrm{J}}$) together in the Sun’s primordial star cluster.
@@
In this paper, we consider close encounters from a typical open star cluster environment using an initial mass function that includes sub-stellar objects with masses ranging between $10^{-3}\,\mathrm{M}_{\odot}$ and $10^{2}\,\mathrm{M}_{\odot}$. We focus on encounters with perihelia less than $20\,\mathrm{AU}$ of the Sun. Due to the closeness of the flybys, we run full $N$-body simulations rather than analytical estimates to accurately resolve the close encounters. In total, we simulate $5\cdot 10^{4}$ flybys with the solar system giant planets initially on circular, coplanar orbits at their current semi-major axes.
@@ 2.1 Stellar Environment
We use a Monte Carlo approach to investigate the impact of various flybys on the solar system by simulating $5\cdot 10^{4}$ flyby objects from an open star cluster environment with a mass density of $100\,\mathrm{M_{\odot}\,pc^{-3}}$. For each flyby, we sample a mass, velocity, and impact parameter along with a random orientation centered on the Sun (Zink et al., 2020).
The initial mass function (IMF) we use to generate the flyby masses $m_{\star}$ is the Chabrier IMF (Chabrier, 2003) which smoothly combines a log-normal IMF for single stars with masses less than one solar mass and the standard power-law IMF (Salpeter, 1955) for masses above one solar mass. We sample masses from the IMF in the range from $10^{-3}\,\mathrm{M}_{\odot}$ to $10^{2}\,\mathrm{M}_{\odot}$.
We note that there is some uncertainty in the IMF on the low-mass end. Nevertheless, IMFs seem to be in reasonable agreement across different environments and star forming regions as long as the conditions are not too extreme (Bastian et al., 2010). Raghu et al. (2024) show that theoretical models of brown dwarf formation are consistent with the observed power-law of the mass probability distribution function (PDF) obtained by microlensing observations by Chabrier and Lenoble (2023). Yet, Kirkpatrick et al. (2024) find that previous IMFs might over-estimated the abundance of some portions of the substellar mass regime.
Because of this uncertainty we compare various popular IMFs (Kroupa, 2001; Chabrier, 2003, 2005; Maschberger, 2013; Kirkpatrick et al., 2024) in Figure 1. Each IMF is normalized assuming a mass density of $100\,\mathrm{M_{\odot}\,pc^{-3}}$.
Figure 1: A comparison of IMFs. While there is uncertainty in the substellar regime, most IMF models are consistent in modeling the number density of substellar objects within one order of magnitude. This paper uses the Chabrier (2003) IMF. We also highlight the approximate mass range for which flybys successfully produce a dynamical architecture similar to the solar system (see also Fig. 2).
For the IMFs shown, the number of objects per cubic parsec, $n$, varies from $181\,\mathrm{pc^{-3}}$ (Chabrier, 2005) to $332\,\mathrm{pc^{-3}}$ (Maschberger, 2013). Roughly half of the objects have substellar masses, $m<80\,\mathrm{M_{J}}$, corresponding to an expected number density of substellar objects ranging from $52\,\mathrm{pc^{-3}}$ to $175\,\mathrm{pc^{-3}}$, respectively.
We sample hyperbolic excess velocities $v_{\infty}$ from a Maxwell-Boltzmann distribution with a velocity dispersion of $1\,\mathrm{km\,s}^{-1}$ (Girard et al., 1989; Adams, 2010). We sample impact parameters $b_{\star}$ to uniformly cover the cross-sectional area out to $75\,\mathrm{AU}$. Gravitational focusing (Binney and Tremaine, 2008) naturally gives a perihelion which is less than the impact parameter with the host star. For our simulations, all perihelia come within $55\,\mathrm{AU}$ with $90\%$ coming within $20\,\mathrm{AU}$. We sample the incident angles isotropically to allow for all possible encounter geometries. We initialize the solar system with the four giant planets on circular and coplanar orbits at their current semi-major axes, and randomly initialize their true anomalies.
@@ 2.3 Metric for secular architectures
For the purpose of comparing the orbital architecture of two systems, some previous studies have used the angular momentum deficit (AMD), an approximately conserved quantity of the solar system planets (Laskar, 1997, 2000; Turrini et al., 2020). Others have used a comparison of the instantaneous eccentricities and inclinations (or their minimum-to-maximum ranges) for each planet (Marzari and Weidenschilling, 2002; Tsiganis et al., 2005; Morbidelli et al., 2007). At a more granular level, some studies have used the amplitudes of the fundamental secular modes for quantitative comparisons of the secular architectures of model solar systems (Morbidelli et al., 2009; Brasser et al., 2009). In the context of planet-planet close encounters in the solar system, Nesvornỳ and Morbidelli (2012) describe a four-step procedure to measure the success of planet-planet interactions accounting for the secular architecture of the solar system. In the Appendix (A.2–A.4), we comment further on these previously used metrics compared to our adopted metric, described below.
For a more comprehensive comparison, we would like to measure the closeness of not only the fundamental secular modes but also their long-term variations, since the fundamental secular modes are merely the linearized approximation of an inherently non-linear system (cf. Lithwick and Wu, 2014). All the above reasons motivate us to develop a new metric which measures the distance of a simulated system from the real solar system without making any initial assumptions regarding their dynamical structure or architecture.
Let us start by defining the root-mean-square log-spectral distance, a measure of the distance between two Fourier spectra, often used in signal processing applications such as speech recognition (Gray and Markel, 1976; Rabiner and Juang, 1993). It compares the power spectra of two signals to evaluate the similarity of the amplitudes for each frequency as
@@
where $e_{j}$ is the eccentricity, $i_{j}$ is the inclination with respect to the invariant plane, $\Omega_{j}$ is the longitude of the ascending node, $\varpi_{j}=\omega_{j}+\Omega_{j}$ is the longitude of the periapsis, $\omega_{j}$ is the argument of the periapsis, and $\mathtt{i}=\sqrt{-1}$. We sample these values on evenly spaced time intervals over 20 Myr integrations with $N=2^{11}$ samples to then process with an FFT. We use $\mathcal{D}$ in Eq. 1 to build a metric that evenly weighs the signals of the complex eccentricities and complex inclinations for each planet. We define our metric as
@@
2.4 Reference Set of Solar Systems
In order to make a fair comparison between the flyby systems and the observed solar system, we use a reference set of solar system realizations from an ensemble of integrations (Brown and Rein, 2020). This ensures that our reference set representing the solar system’s secular architecture samples not just one single 20 Myr realization but also samples the effect of uncertainties of the initial conditions and of the small changes in the solar system’s power spectrum over long times, both of which arise from nonlinear effects.
We start with the nominal initial conditions on 1 January 2000, obtained from NASA’s Jet Propulsion Laboratory (JPL) Horizons data service. Our solar system ensemble uses these nominal initial conditions that are identical in every way except for a small variation ($0.5\,\mathrm{m}$) in the semi-major axis of Mercury from the present-day value (see Brown and Rein 2020 for more details). We use all $96$ of the 5 Gyr solar system simulations and divide them into 20 Myr non-overlapping segments from which we sample the complex eccentricities and inclinations to process with our metric. We note that the fundamental secular mode frequencies of the giant planets change by at most $15\,\mathrm{mas\,yr}^{-1}$ and their amplitudes vary by at most $9.5\cdot 10^{-4}$, but that the power at other frequencies is more variable amongst the 20 Myr snapshots of the solar system taken over billions of years (see Figure 5).
Using the metric $\mathcal{M}$ and this set of 20 Myr segments from our reference solar system ensemble, we cross compare each segment of a random subset of $2500$ segments against another randomly selected subset of $2500$ segments to compute the distribution of values. This distribution of values of $\mathcal{M}$ quantifies how similar the solar system is to itself at different times and with slightly different initial conditions (within observational uncertainties). The range of values given by $\mathcal{M}$ for this cross comparison of the solar system are between 0 and $0.434$.
@@ 2.4 Reference Set of Solar Systems
We found that this procedure provides a clear way of evaluating how close two systems are to each other and, most importantly, allows us to sort and rank them in closeness. Moreover, this procedure avoids over-fitting to a specific manifestation of the solar system’s long and diverse dynamical evolution. What is not immediately clear is how to define the cutoff value of $\mathcal{M}$ that determines if a given system is considered to be a close match to the solar system or not. For the present study, we define any flyby system with a metric value that overlaps with the reference solar system distribution as a close match to the solar system. We acknowledge that this choice is somewhat arbitrary and comes with several caveats. Note that the maximum value of $\mathcal{M}$ could be limited by the ensemble size. Our ensemble size is large, $2500^{2}>6\cdot 10^{6}$, so we don’t expect this to be important in practice. Nevertheless, there exists the rare possibility that a solar system trajectory will eventually go unstable, for example with a Mercury–Venus collision (Laskar et al., 2004; Laskar and Gastineau, 2009). The closer we get to such an instability, the more the trajectory diverges from what is a more ‘typical’ realization of the solar system. Such a divergent trajectory will cause the range of metric values to increase significantly above the typical range and thus lead to more closely matching flyby simulations. However, if there aren’t any significantly divergent trajectories (as is the case with our solar system ensemble, see Brown and Rein 2020), then to use anything less than the entire range of the metric would be to claim that the solar system is not a close match to itself. In summary, although it may be possible to find a better metric, the definition of what is considered a close match to the solar system will likely always be subject to some reasonable but arbitrary quantitative choices.